Various financial data for SunPath Manufacturing for 2015 and 2016 follow.

2015 2016
Output: Sales $300,000 $330,000
Inputs: Labor $40,000 $43,000
Raw Materials: $45,000 $51,000
Energy: $10,000 $9,000
Capital Employed: $250,000 $262,000
Other: $2,000 $6,000

What is the percentage change in the energy partial productivity measure for SunPath between 2015 and 2016?

2.33%

-9.22%

22.2%

-0.53%

2.88%
Answer #1
Energy partial productivity = Output Sales / Energy
Energy partial productivity - 2015= 300,000 / 10,000 = 30
Energy partial productivity -2016 = 330,000 / 9,000 = 36.67
Percentage change in the energy partial productivity = (36.67-30)/30
Percentage change in the energy partial productivity = 22.2%
Hence, the correct option is 22.2%.